ISSN 0193–9025

Food Research Institute Studies

CONTENTS

Deliveries on Chicago Board of Trade

Wheat, Corn, and Soybean Futures Contracts, 1964/65–1988/89 Anne E. Peck and Jeffrey C. Williams

Vol. XXII, No. 2, 1991

FOOD RESEARCH INSTITUTE STUDIES Vol. XXII, No. 2, 1991

Announcing a New Publication by Food Research Institute Scholars and Colleagues

Rice Policy in Indonesia

Scott Pearson, Walter Falcon, Paul Heytens, Eric Monke, and Rosamond Naylor

This book examines the components of Indonesia’s rice policy to assist policymakers, analysts, and observers in sorting out the pros and cons of alternative courses of action. The volume contains the results of the Food Research Institute’s third multi-year research project on Indonesian food policy. It combines new field-based empirical evidence on rice farming profitability and rural employment and wages with the authors’ long experience in analyzing Indonesian food policy issues and the international market for rice.

Rice Policy in Indonesia will be valuable reading for food policy analysts, agricultural economists, agronomists, agricultural engineers, plant breeders, and soil scientists, as well as specialists interested in Indonesia and development studies.

Scott Pearson is Professor and Walter Falcon is Professor and Director at the Food Research Institute, Stanford University. Paul Heytens is Project Economist at the Asian Development Bank, Manila, Philippines. Eric Monke is Professor of Agricultural Economics, University of Arizona. Rosamond Naylor is Research Associate at the Institute for International Studies, Stanford University.

Cornell University Press Ithaca and London

$29.95, cloth

Published by the Food Research Institute, Stanford University, Stanford, California 94305. Reprints of articles in this issue may be purchased from the Institute.

Articles in this journal are abstracted or indexed in Journal of Economic Literature, Agrindex, Bibliography of Agriculture, L’Agronomie Tropicale, Biological Abstracts, Biological and Agricultural Index, Field Crop Abstracts, Nutrition Abstracts and Reviews, Public Affairs Information Service Bulletin, Schrifttum der Agrarwirtschaft, Abstracts on Tropical Agriculture, World Agricultural Economics and Rural Sociology Abstracts, Asian-Pacific Economic Literature, the American Economic Association Index of Economic Articles, and AGRICOLA. Table of contents of each issue is also listed in Current Contents/Agriculture, Biology, and Environmental Sciences.

c_{◯}1991 by the Board of Trustees of the Leland Stanford Junior University Annual subscription, $30.00. Single copy, $10.00.

Food Research Institute Studies

Editor: Walter P. Falcon Associate Editor: William O. Jones Associate Editor: Carl H. Gotsch Assistant Editor: Linda W. Perry

Vol. XXII 1991 No. 2

FOOD RESEARCH INSTITUTE STUDIES Vol. XXII, No. 2, 1991 CONTENTS Deliveries on Chicago Board of Trade Wheat, Corn, and Soybean Futures Contracts, 1964/65–1988/89 by Anne E. Peck and Jeffrey C. Williams

1. Introduction

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129

Concerns About Delivieres of Wheat, Corn, and Soybeans, 130—The Plan of the Study, 133.

2. Historical Levels of Deliveries and the Changing Commercial

Importance of the Delivery Locations . . . . . . . . . . . . . . . . . . . . . 137

Levels of Wheat, Corn, and Soybean Deliveries, 137— Changes in the Position of Chicago and Toledo as Trading Centers, 143—Representativeness of the Deliverable Stocks, 146—Conclusions, 151.

3. Concentration in Cash and Futures Markets

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153

Utilization of the Available Delivery Space, 154— Concentration in the Delivery Space, 156—Concentration of Futures Positions, 158—Price Effects of the High Levels of Concentration Relative to Deliverable Stocks, 164—Conclusions, 170.

.

4. Economic Determinants of the Amount,

Timing, and Location of Deliveries . . . . . . . . . . . . . . . . . . . . . . . . 173

The Delivery Process, 174—The Amount Delivered, 175—Timing of Deliveries During the Delivery Month, 181—The Location of Deliveries, 184—Conclusion, 193.

5. Basis Convergence and Hedging Effectiveness

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195

The Effects of the Timing Option on Basis Convergence, 197—Basis Convergence for the CBOT Contracts, 201—Basis Behavior at Non-Deliverable Locations, 208—Hedging Effectiveness and Multiple Delivery Points, 213.

6. Conclusions and Recommendations

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215

Citations

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221

Appendix: Data Sources

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223

Deliveries, 223—Futures Markets Positions, 223— Prices, 224–Stocks, 225.

STAFF OF THE INSTITUTE

Administration

Walter P. Falcon  Anne E. Peck
Director  Associate Director
Susan A. Maher  Charles C. Milford
Department Administrator  Librarian
Carl H. Gotsch  Linda W. Perry
Associate Editor  Assistpant Editor

William O. Jones Associate Editor

Research Staff

W. Brian Arthur, Professor  Anne E. Peck, Professor
Walter P. Falcon, Professor  Clark W. Reynolds, Professor
Carl H. Gotsch, Professor  Pan A. Yotopoulos, Professor
Timothy E. Josling, Professor  Jeffrey Williams, Associate Professor
Reynaldo Martorell, Professor  Sandra O. Archibald, Assistant Professor
Scott R. Pearson, Professor  Marcel Fafchamps, Assistant Professor

Scott D. Rozelle, Assistant Professor

Visiting Scholars

Ben Crow  Leon A. Mears

Emeritus Professsors

Roger W. Gray  Bruce F. Johnston  William O. Jones

Dudley Kirk